# The Learning Approach to Games

### v2

@@ 4 Reinforcement Learning
${\Upsilon}:\mathscr{M}_{Q}\to{\cal P}(\mathbb{A})\qquad{\Upsilon}(Q)=\delta_{\operatorname*{arg\,max}_{a\in\mathbb{A}}Q(x,a)}$
The learning algorithm $\mathfrak{L}\colon\mathscr{E}\to\mathscr{M}_{Q}$ implemented, for example, via temporal-difference or Monte-Carlo updates, iteratively adjusts the estimation to yield higher values for favorable outcomes, while preserving the time-consistency of the estimation.
We note that the standard RL terminology maps naturally to the learning parameters we have introduced. Learning $\hat{p}$ is typically associated with model-based methods. The horizon $\hat{T}$ is often handled by adjusting the discount factor, though more advanced approaches model it as a stopping time within the options framework. The parameter $\hat{F}$, representing rewards, may be predefined or learned through methods such as inverse reinforcement learning. The function $\hat{\phi}$ plays a role similar to the value function, and its treatment as a random variable aligns with distributional RL. Learning $\hat{\pi}$ corresponds to policy learning, while $\hat{\Gamma}$ is sometimes referred to as behavior prediction. It is important to note that an agent may adopt more sophisticated strategies by learning additional, related parameters. For example, this may include encoding raw observations, establishing communication protocols, modeling opponents’ incentives, or tuning exploration parameters, such as the best expected rewards in the two-player game example in Section 2.1.
To simplify the discussion and notation, we assume that the state and action spaces, $\mathbb{S}$ and $\mathbb{A}$, are fixed and that the estimated functions are time-invariant. In the absence of other players, $\hat{\Gamma}$ becomes irrelevant. The observations ${}^{n}{\cal O}$, which might include states, actions, and rewards, are generated by the environment (or universe) $(\Omega^{u},{\cal F}^{u},\mathbb{P}^{u})$ in which the player operates. The player’s objective is to learn $(\hat{T},\hat{p},\hat{F},\hat{\phi},\hat{\pi})$. Given $(\hat{T},\hat{p},\hat{F},\hat{\phi})$, we have the cost function as
$J(x;\alpha):=\mathbb{E}^{\mathbb{P}^{x,\alpha}}\Big{[}\hat{\phi}(X_{\hat{T}})+\sum_{t\geq 0}^{\hat{T}-1}\hat{F}(X_{t},\alpha(t,X_{t}))\Big{]}$
Then, the player aims to find the optimal controls of $J(x;\alpha)$ using $\hat{\pi}\in{\cal P}({\cal A})$, which yields induced distributions ${\Upsilon}^{x},\gamma^{x}$ as in (2.8), (2.9). We remark that, instead of treating $\hat{F},\hat{\phi}$ as known and then assigning a method of randomizing actions to explore, we model $\hat{F},\hat{\phi}$ as unknown, which then already induces a distribution over actions. This aligns well with the intuition that better characterized values should yield less uncertain strategies.
Let us observe how we can mimic the Markov Decision Processes. Let $(X,{\cal I})$ be the canonical process on $(\mathbb{S}\times\mathbb{A})^{\infty}$. Given $(x_{0},a_{0})\in\mathbb{S}\times\mathbb{A}$, introduce an induced distribution from the model of the player $\hat{p}$ and $\gamma^{x}$ as
{~$\begin{cases}&\tilde{\mathbb{P}}^{x_{0},a_{0}}(X_{s+1}=dy,{\cal I}_{s+1}=d\tilde{a}|X_{s}=x,{\cal I}_{s}=a):=\hat{p}(x,a;dy)\gamma^{y}(d\tilde{a}),\ \ s\geq 0\\ &\tilde{\mathbb{P}}^{x_{0},a_{0}}(X_{0}=x_{0},{\cal I}_{0}=a_{0})=1\end{cases}$
Now, we can introduce a related $Q$-function,
$Q(x,a):=\mathbb{E}^{\tilde{\mathbb{P}}^{x,a}}\Big{[}\sum_{t\geq 0}\lambda^{t}\hat{F}(X_{t},{\cal I}_{t})\Big{]}$
For notation, we didn’t keep track of learning index $n$, but both $\tilde{\mathbb{P}}$ and $\hat{F}$ (and consequently $Q$) are evolving as learning progresses. Note that $Q$ satisfies a Bellman type relation
$Q(x,a)=\hat{F}(x,a)+\lambda\int_{\mathbb{S}\times\mathbb{A}}Q(y,\tilde{a})\ \hat{p}(x,a;dy)\times\gamma^{y}(d\tilde{a})$
Now, if we knew exactly what $p=\hat{p},\ F=\hat{F}$ are, and were able to compute the optimal controls for the infinite horizon, then $\gamma^{x}$ would have support on the maximizers, and $V(x):=\sup_{a}Q(x,a)=\int_{\mathbb{A}}Q(x,\tilde{a})\gamma^{x}(d\tilde{a})$. That is, we recovered the standard Bellman equation:
$V(x)=\sup_{a}\bigg{\{}F(x,a)+\lambda\int_{\mathbb{S}}V(y)\ p(x,a;dy)\bigg{\}}$
As usual, however, it is almost never feasible to know these parameters in reality, nor do they truly exist and fully represent the system at hand. We remark that,
@@
assigns a scalar value for $(x,a)$, which is better suited in our framework as it does not introduce simplifications. We also note that, in simpler settings, it might be more appropriate to learn a single estimate $Q(x,a)$ directly, and avoid all the estimations we have introduced. This is of course a design choice for the player.
We briefly illustrate a method for learning CartPole [21]. The goal is not to propose a novel learning method, but to further emphasize the connections by learning $\hat{\phi}$ under a fixed horizon $T=8$. Let us revisit the basics of the CartPole problem. The state space is $\mathbb{S}=\mathbb{R}^{4}$ representing position, velocity, angle, and angular velocity. The action space is $\mathbb{A}=\{0,1\}$, where the actions represent applying force to the left or right of the cart. The player gains a reward at each time step if the pole is vertical within some bounds.
The player has a memory for recording observations, including states, actions, and episode-level evaluations. After each episode, a performance metric is computed as the average of the agent’s relative and absolute performance.
In this problem, $\hat{p}$ is deterministic and can be learned. However, as it is not of our interest, we took it as given. We also set $\hat{F}=0$ to only model the state values $\hat{\phi}$. Introduce $\{{\cal N}_{\phi}^{k}\}_{k=1}^{K_{\phi}}$ for some $K_{\phi}\in\mathbb{N}$ as neural networks $\mathbb{S}\to\mathbb{R}_{+}$ for state values. Let $\mathfrak{L}_{\phi}:\mathscr{E}\to(\hat{\Omega}\times\mathbb{S}\to\mathbb{R}_{+})$
$\mathfrak{L}_{\phi}({}^{n}{\cal O}):=\sum_{k=1}^{K_{\phi}}\delta_{{\cal N}_{\phi}^{k}}{\bf 1}_{\left\{{k}\right\}}(\hat{\omega}),\quad\hat{\omega}\in\hat{\Omega}:=\{1,\dots,K_{\phi}\},\ \text{ and }\ \hat{\mathbb{P}}\text{ uniform}$
Here, $\hat{\mathbb{P}}$ taken as uniform means that the player has no information about which network is providing better estimations, which is again for simplicity only. We took the range of each ${\cal N}_{\phi}^{k}$ as $[0,100]$, and trained these networks relying on the memory. The basic idea is that networks are trained to map the states to their corresponding performance values. Here, we randomly sample from best and worst performances in a balanced way, creating a diversity between ${\cal N}_{\phi}^{k}$’s. Moreover, we also trained networks to promote time-consistency.
To ensure that the value networks are properly trained, we consider the set of all controls $\mathbb{A}^{T}=\{0,1\}^{T}$ and index them by $\{\alpha^{k}\}_{k=1}^{2^{T}}$. [^12] We note that it is considerably easier learning controls in this simple setting, since $\mathbb{A}^{T}$ is finite. Once we include policy learning, the player performs quite well even without properly trained value networks, and hence omitted. Then, we let $\mathfrak{L}_{\pi}:\mathscr{E}\to(\hat{\Omega}\times\mathbb{S}\to{\cal P}(\mathbb{A}^{T}))$ as
$\mathfrak{L}_{\pi}:=\frac{1}{Z}\sum_{k=1}^{2^{T}}\exp\big{(}J(\hat{\omega},x,\alpha^{k})\big{)}\delta_{\alpha^{k}},\qquad(\hat{\omega},x)\in(\hat{\Omega},\mathbb{S})$
$\text{where}\ \ J(x;\alpha):=\mathbb{E}^{\mathbb{P}^{x,\alpha}}\big{[}\hat{\phi}(X_{T})\big{]}$
and $Z$ is the normalizing constant. Lastly, given $(\hat{\phi},\hat{\pi})$ as $(\mathfrak{L}_{\phi},\mathfrak{L}_{\pi})({}^{n}{\cal O})$, the induced distribution in (2.8) is simply
${\Upsilon}^{x}$ $:=\int_{\hat{\Omega}}\frac{1}{Z}\sum_{k=1}^{2^{T}}\exp\big{(}J(\hat{\omega},x,\alpha^{k})\big{)}\delta_{\alpha^{k}}d\hat{\mathbb{P}}(\hat{\omega})=\frac{1}{K_{\phi}Z}\sum_{\ell,k=1}^{K_{\phi},2^{T}}\exp\big{(}J(\ell,x,\alpha^{k})\big{)}\delta_{\alpha^{k}}$
Figure 2: Performance of the CartPole Game Across 10 Selected Runs. The x-axis represents the number of games (or episodes), while the y-axis shows the total reward for each episode. Each line corresponds to one of the best 10 runs out of 32, with small markers indicating individual episode scores and solid lines showing the moving average calculated from up to the last 100 episodes.
Lastly, we will briefly examine the simplest case of Bandit Games and demonstrate how it can be embedded into our framework in two different ways. The first approach adopts a single-player perspective, where the arms are represented by the state $X$. The second approach treats arms as players who take actions but do not engage in learning. Both versions will help to explore our framework further, and may generalize in different directions.
In the state perspective, let $\mathbb{S}=(S^{1},\dots,S^{k})$ represent a fixed $k$ dimensional state space. Here, transition probabilities are defined by $(\mu^{1},\dots,\mu^{k})$, which are independent of the current state, time and actions, representing the simplest dynamics of each arm. Let us note that random samplings from these distributions characterize $\mathbb{P}^{u}$, the universal probability distribution. In this perspective, the state of the chosen arm corresponds to the reward of the main agent, and the learning objective is $\hat{p}$ aiming to learn $(\mu^{1},\dots,\mu^{k})$.
Fix $\hat{T}=1$, and $\mathbb{A}=\{1,\dots,k\}$. As the reward of the agent depends on the current action and the next state, introduce
$J(\hat{\omega},a):=\mathbb{E}^{\mathbb{P}}\big{[}\phi(a,X_{1})\big{]}+\hat{F}(\hat{\omega},a),$
@@
Note that, we adapted the framework in a way $\phi$ does not necessarily corresponds to the standard value function, and $\hat{F}$ aims to capture the underlying distribution of the state.
It is important to note that an adversary can be included as another player within the model. The adversary might observe the actions of the agent, and the adversary’s actions can alter the underlying transition probabilities of the states. This introduces potential extensions, such as scenarios where both the agent and the adversary are attempting to estimate each other’s actions.
In the action perspective, let players $1,2,\dots,k$ represent arms in the game. Each arm has a fixed horizon $\hat{T}=1$, and transition probabilities or estimating others’ actions are not relevant for the arms. We characterize the reward of the main agent as the action taken by the chosen arm. For the rewards of the arms, one can easily define $\hat{F}^{i}(\hat{\omega},a)$ on a probability space $(\hat{\Omega},\hat{\mathbb{P}})$ such that the induced distributions over actions correspond to the targeted reward distribution $\mu^{i}$.
Here, the objective of the main agent is to characterize $\hat{\Gamma}$ as defined in (2.4). Let $\hat{\Gamma}$ be independent of the agent’s actions, $\hat{\Gamma}\in{\cal P}(\mathbb{A}^{k})$, where $\mathbb{A}$ represents the action space of the arms (which corresponds to the rewards of the agent). Consequently, we can similarly define the cost function for the main agent as
[…]
The cost function is exactly the same as in the state perspective; however, it is represented as an expectation over $\hat{\Gamma}$ rather than $\hat{p}$.
In this perspective, it is more natural to consider each arm as an individual adversary. These adversaries can cooperate, as cooperation might be manifested through $\hat{\Gamma}^{i}$s that estimate the actions of other adversaries. Furthermore, the main agent might aim to capture the correlated behaviors of the arms.
We stress that we have not designed any learning algorithm that constitutes an uncertain equilibrium, nor have we characterized its stability, regret bounds, or other related metrics. In the simplest scenario, where there is one optimal arm and the reward distributions are fixed, any reasonable algorithm that successfully identifies this arm would correspond to an uncertain equilibrium. However, in more complex and realistic scenarios, the specific setup must be further explored to develop a favorable learning algorithm, which falls outside the scope of this work.

### v4

@@ 5 Reinforcement Learning
${\Upsilon}:\mathscr{M}_{Q}\to\mathscr{M}_{\Upsilon}\qquad{\Upsilon}(Q)=\operatorname*{arg\,max}_{a\in\mathbb{A}}Q(\cdot,a)\in\{\mathbb{S}\to\mathbb{A}\}$
The learning algorithm, $\mathfrak{L}_{Q}\colon\mathscr{E}\times\mathscr{M}_{\varphi}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{Q}$, implemented, for example, via temporal-difference updates, iteratively adjusts the estimates to assign higher values to favorable outcomes while preserving their time consistency. In this method, enforcing consistency before the value estimates are well trained often leads to instability. To mitigate this, one may introduce a slower-moving “target” network to anchor the estimates while continuing to explore, employ two value networks with decoupled or conservative targets to reduce overestimation bias or apply additional regularization terms to stabilize both the estimation and behavioral updates. See, for example, [55, 87, 86, 28, 31] for foundational and modern off-policy implementations.
In on-policy learning, the estimates aim to capture the value of the current behavior. This mitigates instabilities that arise from enforcing consistency early in training, but at the expense of making past observations less useful because the behavior continually evolves. See, for example, [91, 81, 54]. For training such estimates, which we do not elaborate on here, see also [70, 69] for important methodological considerations.
Distributional RL addresses the setting in which estimates represent distributions rather than expectations, with estimates taking the form $\mathscr{M}_{\varphi}:\{\mathbb{S}\times\mathbb{A}\to{\cal P}(\mathbb{R})\}$. See, for example, [8, 22, 21, 93]. In the next subsection, we present a toy example in which the estimates are random variables, $\mathscr{M}_{\varphi}:\{\hat{\Omega}\times\mathbb{S}\to\mathbb{R}\}$, similar to the concept of ensemble methods such as [60]
Model-based RL incorporates further spaces of estimations for the upcoming observations. As we briefly mentioned in Section 2, such estimates may include future states, rewards, and actions. Furthermore, one might first consider embeddings of the observations to facilitate predicting future embeddings. See, for example, [30, 33, 32] for some modern implementations.
Once players have an estimate for future observations, they can plan for the future in a time-inconsistent manner. Similar to the consideration in Section 3, suppose an estimated policy takes values in the space of controls rather than actions. In such cases, the plan devised for a potential future observation may differ from what is actually planned when that situation is realized. This is expected, as the player tailors its efforts to the current situation. This is an important generalization, which is typically absent in model-free approaches, and suggests that time-consistency is an important but not fundamental property. We will mention three approaches in this direction:
$\bullet$ Model Predictive Control and its modern applications in RL can be seen as a direct example. See, for instance, [57, 20].
$\bullet$ Monte Carlo Tree Search methods also plan into the future, using policy solely to better evaluate the current possible actions; see, for example, [74, 75, 68].
$\bullet$ Hierarchical methods include additional estimations that assign goals to the behavior. This enables the generation of diverse strategies and, in the case of model-free implementations, introduces a form of time inconsistency. However, rather than discarding the future plan, the player commits to the assigned goal. This provides the flexibility to behave differently later, depending on the previously assigned goal; see, for example, [3, 56, 48].
Next, we briefly mention methods for learning representations from observations. The main objective is to reduce the complexity of the observations while retaining a representation that is sufficiently rich for the task at hand. Broadly speaking, one may consider predictive or contrastive objectives. For predictive objectives, see for example, [30, 32, 71]. Contrastive objectives aim to leverage invariance and discriminability, grouping ”similar” observations together while separating ”dissimilar” ones. Examples include [85, 46, 78].
Not only learning observations is important, but exploring the diversity of those observations is also crucial. A widely adopted strategy is to provide intrinsic rewards. For example, a player may favor observations where internal estimations are failing to accurately predict their targets. See, for example, [62, 14, 29]. Another approach is to promote unfamiliar states to the player, such as, [5, 4, 67].
To conclude this brief overview, we point out that even when there is only a single decision-maker, setting up an environment through a Markov decision process (or one of its variants) does not capture the full picture. Optimization is, of course, crucial for formalizing which observations are preferred, and it provides strong guidance for designing the players, as previously discussed. However, when many players interact, seeking compatible strategies without accounting for their internal estimations does not necessarily provide clear guidance on which observations are going to be favored by the players. We remark that, in Definition 7 of uncertain equilibrium, each player optimizes their own objective.
5.1 A learning algorithm for CartPole
We first briefly illustrate a method for learning CartPole. The goal is not to propose a better learning method, but a novel one to draw an attention to potential variety of learning algorithms for forming estimations. We learn $\hat{\phi}$ with fixed horizon $T=8$. It is a value-based method with model given. We trained value networks without Bellman-type updates and instead promote such time-consistency afterwards.
Let us revisit the basics of the CartPole problem. The state space is $\mathbb{S}=\mathbb{R}^{4}$ representing position, velocity, angle, and angular velocity. The action space is $\mathbb{A}=\{0,1\}$, where the actions represent applying force to the left or right of the cart. The goal is to keep the pole attached to the cart in the upright position.
The player has a memory, a function of observations, for recording observations, including states, actions, and episode-level evaluations. After each episode, a performance metric is computed as the average of the agent’s relative and absolute performance, and assigned as the evaluation of the episode. Here, the relative performance is computed from the players moving average of how many steps pole was upright. The absolute performance is computed depending on the maximum potential steps, which is $500$.
In this problem, $\hat{p}$ is deterministic and can be learned. However, as it is not of our interest, we took it as given. We also set $\hat{F}=0$ to only model the state values $\hat{\phi}$. Introduce $\{{\cal N}_{\phi}^{k}\}_{k=1}^{K_{\phi}}$ for some $K_{\phi}\in\mathbb{N}$ as neural networks $\mathbb{S}\to\mathbb{R}_{+}$ for state values. Let $\mathfrak{L}_{\phi}:\mathscr{E}\times\mathscr{M}_{\phi}\to(\hat{\Omega}\times\mathbb{S}\to\mathbb{R}_{+})$ and
$\mathfrak{L}_{\phi}({}^{n}{\cal O},{}^{n-1}\hat{\phi}):=\sum_{k=1}^{K_{\phi}}{\cal N}_{\phi}^{k}\ {\bf 1}_{\left\{{k}\right\}}(\hat{\omega}),\quad\hat{\omega}\in\hat{\Omega}:=\{1,\dots,K_{\phi}\},\ \text{ and }\ \hat{\mathbb{P}}\text{ uniform}.$
Here, $^{n}\hat{\phi}:=\mathfrak{L}_{\phi}({}^{n}{\cal O},{}^{n-1}\hat{\phi})$ is representing abstractly the whole training process. $\hat{\mathbb{P}}$ is taken as uniform means that the player has no information about which network is providing better estimations, which is again for simplicity only. We took the range of each ${\cal N}_{\phi}^{k}$ as $[0,100]$, and trained these networks relying on the memory. After $6$ episodes, the player goes over the memory to recall best and worst performances, and forms clusters of states to assign higher or lower values to them. We note that, values are changing as the performance increase over time. Also, the training is not done by assigning expected values, but instead directly assigns higher or lower values. Then, we separately promoted time-consistency of these assignments.
To ensure that the value networks are properly trained, we consider the set of all controls $\mathbb{A}^{T}=\{0,1\}^{T}$ and index them by $\{\alpha^{k}\}_{k=1}^{2^{T}}$. [^10] We note that it is considerably easier learning controls in this simple setting, since $\mathbb{A}^{T}$ is finite. We first included a separate policy learning to generate a small amount of controls, and aimed to use value networks to select from them, but the player performs quite well even without properly trained value networks. Hence, we omitted from the discussion and evaluate all of the potential controls to make sure value is well-trained. Then, we let $\mathfrak{L}_{\pi}:\mathscr{M}_{\phi}\to(\tilde{\Omega}\times\hat{\Omega}\times\mathbb{S}\to\mathbb{A}^{T})$ as
$\mathfrak{L}_{\pi}=\sum_{k=1}^{K_{\phi}}\sum_{\ell=1}^{2^{T}}\delta_{\alpha^{\ell}}{\bf 1}_{\left\{{l}\right\}}(\tilde{\omega}){\bf 1}_{\left\{{k}\right\}}(\hat{\omega})\quad\text{where,}\ \ \tilde{\omega}\in\tilde{\Omega}=\{1,\dots,2^{T}\},\ \$
$\hat{\mathbb{P}}(\tilde{\omega},\hat{\omega})=\frac{1}{Z}\exp\big(J(\hat{\omega},x,\alpha^{\tilde{\omega}})\big),\ \ \text{and}\ \ J(\cdot,x;\alpha)=\mathbb{E}^{\mathbb{P}^{x,\alpha}}\big[\hat{\phi}(\cdot,X_{T})\big]$
Here, we extended the probability space by adding $\tilde{\Omega}$ to separately keep track of randomness coming from value networks, and the policy that they induce. $\hat{\mathbb{P}}$ is also now viewed as a probability over $\tilde{\Omega}\times\hat{\Omega}$, and $Z$ is the normalizing constant.
Figure 2: Performance of the CartPole Game Across 10 Selected Runs. The x-axis represents the number of games (or episodes), while the y-axis shows the total reward for each episode. Each line corresponds to one of the best 10 runs out of 32, with lines showing the moving average calculated from up to the last 100 episodes.
@@ 5.1 A learning algorithm for CartPole
In fact, the behavior commits for $T$ steps, but that would need to introduce the parameter $t$ and we omitted it for brevity. See Figure 2 for the performance, and refer to the repository [41] for implementation details. Automatically generated reports, produced using LLMs, are also provided to enhance the accessibility of the implementation.
5.2 Multi-agent Reinforcement Learning
Our aim in this section is to provide a broad view of the multi-agent reinforcement learning literature through the lens of the language developed in this work. Once multiple players are present, relevant considerations expand dramatically. Agents may model opponents, communicate, teach, manipulate, or form temporary alliances, and it is difficult to place principled bounds on such possibilities. In particular, if a collection of agents is capable of developing a language to communicate observations or situations they have encountered, this opens the door to higher-order interactions: agents may revisit past experiences, reassess estimations, propose alternative plans, negotiate over future actions, implement voting mechanisms, or assign specialized roles within the group. While such behaviors are routine for us, they illustrate how rich multi-agent learning can become once agents are viewed as structured decision-makers rather than as static policies.
We begin by mentioning a major line of work in multi-agent learning, which is motivated by convergence to equilibrium via regret minimization. Initiated by [95] in games with incomplete information, these methods iteratively update policies so as to minimize regret, thereby converging to unexploitable strategies. This approach has led to notable success in heads-up limit Hold’em poker [10]. Subsequent work has focused on scaling beyond tabular representations, including neural network approximations [13] and their integration with search-based planning [12]. This line of research is fundamentally equilibrium-oriented. From the perspective of a central designer seeking stable strategies to announce, such methods play a crucial role. However, as our focus is on player design rather than equilibrium computation, we do not pursue this direction further. We refer the interested reader to earlier equilibrium-motivated approaches based on Q-learning [38], and to [6] for an analysis of convergence dynamics in differentiable games.
To extend value-based designs to multi-agent settings, one direct approach is to treat each player as unaware of the others and to regard the remaining players as an unknown component of the environment. In this case, each player $({\cal O}^{i},\mathfrak{L}_{Q}^{i},{\Upsilon}^{i})$ still relies on a single type of estimate, such as $\mathscr{M}_{Q}^{i}:=\{\mathbb{S}\times\mathbb{A}^{i}\to\mathbb{R}\}$, together with a common behavior ${\Upsilon}^{i}$. However, unlike the single-agent setting, the optimization target of the learning algorithms is no longer fixed by the environment, since the behaviors of other players are themselves evolving. [^11] For simplicity, we write estimates as functions of the state $\mathbb{S}$. In practice, observations are often processed through additional estimates, such as recurrent neural networks, and the output space of those estimates is used instead. We omit such considerations to focus on what is new in the multi-agent case.
Without allowing players to explicitly acquire additional estimates about their opponents, one can still couple them through a shared estimate. A large body of the literature studies this centralized-training–decentralized-execution regime, primarily in cooperative settings. Since the learning algorithm is centralized, the training phase is more naturally viewed as a single player $(\vec{\cal O},\mathfrak{L}_{Q},\vec{\Upsilon})$ with high-dimensional observations and behaviors.
Consider, for example, a shared estimate space $\mathscr{M}_{Q}:=\{\mathbb{S}\times\vec{\mathbb{A}}\to\mathbb{R}\}$ with learning algorithm $\mathfrak{L}_{Q}:\vec{\mathscr{E}}\times\vec{\mathscr{M}}_{Q}\times\mathscr{M}_{Q}\to\vec{\mathscr{M}}_{Q}\times\mathscr{M}_{Q}$. As long as ${\Upsilon}^{i}$ depends only on $\mathscr{M}_{Q}^{i}$, once training is complete, one can define individual players as $({\cal O}^{i},\mathfrak{L}^{i},{\Upsilon}^{i})$, where $\mathfrak{L}^{i}$ yields $Q^{i}$ as a constant in order to freeze training.
There are several ways to realize such a single-player formulation. For instance, in [80], the authors define $Q$ as the sum of the individual $Q^{i}$’s. In [65], this approach is generalized by introducing an additional estimate $\varphi:\vec{\mathscr{M}}_{Q}\to\mathscr{M}_{Q}$ satisfying $\partial_{Q^{i}}\varphi\geq 0$, ensuring that the learned estimates $Q^{i}$ respect each player’s local maximization behavior. In [77], the authors further generalize this framework by relaxing the monotonicity constraint and introducing consistency conditions that allow a broader class of decompositions while preserving individual players’ behavior. Finally, in [90], the analysis is carried out in the advantage space by introducing two estimates $V$ and $A$ to define $Q$, while maintaining a general representation.
The same centralization idea can also be applied to actor–critic designs. In this case, the planned behavior $^{n}\vec{\Upsilon}\in\vec{\mathscr{M}}_{\Upsilon}$ of the central (single) agent is fixed after training, and the individual agents are equipped with mappings ${\Upsilon}^{i}$ that constantly yield $^{n}{\Upsilon}^{i}\in\mathscr{M}_{\Upsilon}^{i}$.
In [49], the authors apply a deterministic policy gradient method with a central player $(\vec{\cal O},\vec{\mathfrak{L}}_{Q},\vec{\Upsilon})$, where
$\mathfrak{L}_{Q}^{i}:\vec{\mathscr{E}}\times\vec{\mathscr{M}}_{Q}\times\vec{\Upsilon}\to\mathscr{M}_{Q}^{i},$
${\Upsilon}^{i}:\mathscr{E}^{i}\times\mathscr{M}_{Q}^{i}\times\mathscr{M}_{\Upsilon}^{i}\to\mathscr{M}_{\Upsilon}^{i},$
and $\mathscr{M}_{\Upsilon}^{i}=\{\mathbb{S}^{i}\to\mathbb{A}^{i}\}$, $\mathscr{M}_{Q}^{i}=\{\vec{\mathbb{S}}\times\vec{\mathbb{A}}\to\mathbb{R}\}$. In [25], the authors consider stochastic policy gradients, and hence $\mathscr{M}_{\Upsilon}^{i}=\{\Omega^{u}\times\mathbb{S}^{i}\to\mathbb{A}^{i}\}$ when written as a random variable. They introduce an advantage estimate that compares the current $Q$ estimate to an average taken over the planned behavior $^{n}{\Upsilon}^{i}$, in order to provide individual credit assignment. In [94], the authors demonstrate the effectiveness of proximal policy optimization in this regime. This core method has also been successfully used in large-scale systems such as OpenAI Five for Dota 2 [59].
In this centralized regime, communication has also been studied extensively in the literature. In this context, communication often refers to the introduction of two additional estimates. First, each agent constructs messages from its individual observations, taking values in a Euclidean space, $C\in\mathscr{M}_{C}^{i}:=\{\mathscr{E}^{i}\to\mathbb{R}^{k}\}$. Second, each player maintains an estimate that aggregates the incoming messages, $D\in\mathscr{M}_{D}^{i}:=\{\vec{\mathbb{R}}^{k}\to\mathbb{R}^{k^{\prime}}\}$. Within this framework, the learning algorithm of the central player can, for example, be formulated as
$\mathfrak{L}_{Q}:\vec{\mathscr{E}}\times\vec{\mathscr{M}}_{C}\times\vec{\mathscr{M}}_{D}\times\vec{\mathscr{M}}_{Q}\times\mathscr{M}_{Q}\to\vec{\mathscr{M}}_{C}\times\vec{\mathscr{M}}_{D}\times\vec{\mathscr{M}}_{Q}\times\mathscr{M}_{Q}$
allowing the estimates $C$ and $D$ to be updated jointly with the value estimates, for instance via backpropagation. The value functions may then incorporate the aggregated messages in their domain, $\mathscr{M}_{Q}^{i}:=\{\mathbb{S}\times\mathbb{A}^{i}\times\mathbb{R}^{k^{\prime}}\to\mathbb{R}\}$, or, alternatively, the planned behavior itself may be defined to depend explicitly on the aggregated messages.
In [26], the authors implement communication both in independent Q-learning and in a centralized variant, following the general structure described above. In [23], the authors augment messages with a notion of directionality, allowing each agent to attend to incoming messages along a chosen direction. In [79], the authors adopt a centralized single-agent perspective, and introduce multiple communication layers $D^{1},D^{2},D^{3},\dots$, in which messages are iteratively averaged. This architecture enables agents to communicate at higher levels, even when they are not directly connected. In [76], the authors extend this line of work to settings with individual agents, including competitive scenarios, and introduce a gating mechanism that allows agents to learn when to exchange messages and when to remain silent. In [44], attention mechanisms are used to aggregate messages across multiple communication layers. Finally, in [66], the attention-based framework is further extended by incorporating a gating mechanism through a learnable adjacency matrix, treating the communication graph itself as a differentiable estimate.
A large-scale project worth mentioning in the context of communication is [53], which studies the turn-based game Diplomacy, a setting that allows natural-language communication between players. The authors train a large language model on human gameplay data to generate dialogue and introduce explicit estimates of opponents’ intents. Policies are then learned in a manner grounded in human play, integrating strategic decision-making with models of communication and inferred opponent intentions. In effect, this approach constructs a richly structured model of a human player. This example further illustrates that such games cannot be adequately captured by standard external formulations.
We remark that the above example already includes elements of opponent modeling. Let us mention a few important works on opponent modeling. Depending on the structure of the game, it may suffice for a player to observe how other agents influence the state of the game. For example, in another large-scale project [88], the authors train independent agents for the game StarCraft II, where players compete through carefully designed self-play as well as against opponents constructed to exploit their strategies, with policies guided in part by human gameplay data. However, a more sophisticated player may need to maintain explicit estimates dedicated to its opponents. In [36], the authors introduce an additional estimate that constructs an embedding of the opponent based on observations. First, they extend the $Q$-function to incorporate this embedding as an additional input. Second, they consider a collection of $Q$-functions and use the learned embedding to assign weights to them. In these approaches, the learning algorithm for the opponent embedding is initially embedded within the learning procedure for the standard $Q$-estimate. Finally, the authors propose introducing a separate learning algorithm specifically for the opponent embedding in order to enhance its expressiveness. In [64], rather than constructing a full model of a player, the authors focus on learning a particular estimate for characterizing an observed opponent. They introduce one estimate that aggregates information across observations from multiple episodes of the observed player, and a second estimate that depends on the first while incorporating partial observations within a single episode. Together with the current observation, the outputs of these two estimates are then used as inputs to a third estimate, whose learning objective is to predict properties of the opponent being observed. In [43], the authors introduce intrinsic rewards for influencing the behavior of other agents, and demonstrate that accounting for such influence can promote coordination and communication. They further show that these intrinsic rewards can be trained in a decentralized manner, provided that agents acquire predictive estimates of other agents’ action distributions.
We have aimed to emphasize the importance of player design. The literature is extensive and spans many different directions. To illustrate how broad the design space can be, we conclude by pointing to recent works in which players incorporate large language models as part of their estimates. For example, [61] demonstrates the emergence of believable human behavior, while [63, 37] introduce teams of agents designed for collaborative software development.
6 Conclusion
We have reframed the problem of decision-making from the perspective of the player and, in essence, abstracted the constructions in reinforcement learning. Traditional approaches adopt a viewpoint external to the players. Even when there is a single decision-maker, the external setting may guide the design but does not suffice to capture the required complexity. When many players interact, considering only compatible strategies is not a sufficiently rich objective, particularly for competing players.
Our broader vision is to emphasize the inherent complexity of intelligent behavior. We do not act through a single estimate, but through dynamically evolving layers of estimations and behaviors: structures capable of working with any stream of observations, adapting to novel situations, reconstructing, planning, and predicting; forming diverse values both individually and collectively. Understanding how a dynamical system achieves such complexity remains far beyond our reach, but this work aims to provide the foundational definitions for initiating a formal approach.